OpenStudy (anonymous):

Solve the system by using a table. {-0.2x -0.3y = 0.2 {0.4x -0.3y = 1.4

OpenStudy (anonymous):

0.4( -0.2x -0.3y = 0.2) 0.2(0.4x -0.3y = 1.4) -0.8x -0.6y = 0.8 0.8x -0.6y=2.8 -1.2y=3.6 y=-3 take any equation and put the y value on it, and you could find the x